OpenStudy (anonymous):

Which of the following is a polynomial function in standard form with zeros at -8, -1, and 3? f(x) = x^3 + 6x^2 - 19x - 24 f(x) = (x - 8)*(x - 1)*(x + 3) f(x) = (x + 8)*(x + 1)*(x - 3) f(x) = x^3 - 6x^2 - 19x + 24 I *think* it's C because that's adding the numbers there listed to the ones above would equal zero but I'm not sure

OpenStudy (anonymous):

yes you are correct factoring is sometimes called finding the zeros so when you completly factor a polynomial is is also called finding the zeros

OpenStudy (anonymous):

however c is in factored form not standard form

OpenStudy (anonymous):

i would factor a and d and see which one factors to the desired zeros

OpenStudy (anonymous):

How do I factor the polynomials?

OpenStudy (paxpolaris):

expand C instead ... and see if you get A or D

OpenStudy (paxpolaris):

\[f(x)\\=(x+8)\left[ (x+1)(x-3) \right]\\=\left( x+8 \right)\left[ x^2-2x-3 \right]\\=?\]

OpenStudy (anonymous):

i was just about to tell him to multiply all the factors together

OpenStudy (anonymous):

I got -24...?

OpenStudy (anonymous):

multiply (x+8)(x+1)(x-3) you should not get -24

OpenStudy (anonymous):

pax polaris got it started

OpenStudy (anonymous):

(x+1)(x-3)= x^2-2x-3

OpenStudy (anonymous):

so multiply (x^2-2x-3)(x+8)

OpenStudy (anonymous):

-36....not quite sure if I'm doing this right

OpenStudy (anonymous):